Nuts to You Shoppe sells cashews for $15 per kg and pecans for $10 per kg. How many kilograms of each should be mixed in order to get 20 kg of a mixture worth $12 per kg.
Answer by mananth(16946)

Price $ ---------------- quantity
Cashews 15 ---------------- x kg
pecans 10 ---------------- 20 - x kg
Mixture 12.00% ---------------- 20

15x+10(20-x)=12 *20

15x+200-10x=240
15x-10x =240-200
5x=40
/5
x=8 kg $15.00 Cashews
12 kg $10.00 pecans